# Scalability Analysis of 5G-TSN Applications in Indoor Factory Settings

### v1

@@ Abstract
While technologies such as Time-Sensitive Networking (TSN) improve deterministic behaviour, real-time functionality, and robustness of Ethernet, there is a drive for future industrial networks to be increasingly wireless. While wireless networks facilitate mobility, reduce cost, and simplify deployment, they do not always provide stringent latency constraints and highly dependable data transmission as required by many manufacturing systems. The advent of 5G, with its Ultra-Reliable Low-Latency Communication (URLLC) capabilities, offers potential for wireless industrial networks. 5G guarantees elevated data throughput, very low latency, and negligible jitter. As 5G networks typically include wired connections from the base station to the core network, integration of 5G with time-sensitive networking is essential to provide rigorous QoS standards. The paper assesses the scalability of 5G-TSN for various indoor factory applications and conditions using OMNET++ simulation. Our research shows that 5G-TSN has the potential to provide bounded delay for latency-sensitive applications in scalable indoor factory settings.
@@ I Introduction
With the rise of the Industrial Internet of Things (IIoT) and the push towards Industry 4.0, efforts are being made to enhance the flexibility, interoperability, and integration of industrial automation systems. The future of smart manufacturing will see the emergence of dynamic, reconfigurable production systems capable of rapid adaptation to evolving industry requirements. These environments will leverage technologies such as mobile sensing devices, robotic arms, automated guided vehicles (AGVs), and drones to facilitate the customization of products. Since traditional wired automation systems lack the adaptability needed for such changes, wireless communication technologies are becoming increasingly crucial to ensure real-time, deterministic connectivity required for critical industrial processes.
Wireless technologies are advancing to support the demanding needs of industrial environments. Wireless TSN, building on the success of wired TSN, is emerging as a promising solution [1]. 5G and 802.11 are potential candidates for enabling wireless TSN. However, 5G, with its ultra-reliable, low-latency communication capabilities, specifically designed for real-time communication through its URLLC profile, appears to be poised to become the wireless backbone for Industry 4.0 applications.
5G and TSN integration promises to significantly advance real-time industrial communication. However, challenges such as precise time synchronization and efficient resource allocation remain. While early testbeds to study 5G-TSN performance are still under construction, further research is also necessary to create tools to assess the scalability of 5G-TSN in industrial settings.
While there are several studies on 5G-TSN systems, to the best of our knowledge, this is the first study to examine 3GPP standardised 5G-TSN in a wireless manufacturing environment and analyse its scalability. The conducted simulation-based study provides a performance assessment of an industrial use case by integrating an indoor manufacturing model into a 5G-TSN industrial network model transmitting TSN traffic over 5G. Our research presents the following key contributions:
@@ II-A4 Resource management
Currently, 5G is recognized as the leading candidate for implementing wireless TSN solutions. In its Release 16, the 3rd Generation Partnership Project (3GPP) laid a strong groundwork for enabling mobile and wireless communications to achieve the speed and reliability required by Industry 4.0. This release introduces mechanisms to deliver ultra-reliable low-latency communication (URLLC) with sub-1 ms latency and highly precise synchronization, with an accuracy better than 1 $\mu$s, leveraging IEEE 802.1AS. Integration with IEEE 802.1 TSN is facilitated through the 5G Bridge framework, which incorporates critical elements such as the device-side TSN Translator (DS-TT), user equipment (UE), the radio access network (RAN), and the core network (5GC). The core network architecture is divided into user plane functions (UPFs), featuring the network-side TSN Translator (NW-TT), and control plane functions that include components for managing access, mobility (AMF), sessions (SMF), policies (PCF), and unified data (UDM) [2]. It is essential to understand the influence of wireless channels when incorporating TSN into 5G networks to ensure the effective management of TSN performance requirements. The next section delves into the unique characteristics of indoor factory environments and explores the use-case of 5G-TSN within this context.
@@ III-B Scenario/Use-Case
Automated Guided Vehicles (AGVs) are adaptable platforms that utilize 5G connectivity to support diverse applications. Initially designed for material handling in factory logistics, AGVs and Autonomous Mobile Robots (AMRs) have recently gained prominence in flexible manufacturing and collaborative robotics, taking over tasks traditionally handled by humans. With 5G integration, AGVs enhance operational efficiency by autonomously inspecting factories for hazards and supporting various safety and monitoring functions. In indoor factory environments, deploying multiple AGVs is often practical to efficiently handle a variety of tasks and ensure seamless operations. Their adoption in industrial logistics, particularly in indoor settings, has seen substantial investment and interest from major companies like Amazon, DHL, Foxconn, Tesla, and Samsung.
While practical implementations exist, researchers continue to explore the challenges, scalability, and implementation strategies for 5G-powered AGV/AMR systems [4]. A practical example involves a mobile industrial robot (MiR) equipped with 5G connectivity, featuring a battery, LiDAR-enabled digital camera, video processing unit, and 5G radio access unit connected to a private 5G network. The setup integrates mobile edge computing for remote access and enhanced functionality, as shown in Fig. 2.
In our 5G-TSN network scenario, traffic is categorized into three types. Network Control (NC), Video - to partly support navigation, and Best Effort (BE). NC and Video are TSN traffic streams, with NC assigned the highest priority, Video a medium priority, and BE, a non-TSN stream, the lowest priority. NC traffic is periodic, with intervals ranging from 50 ms to 1 second and packet sizes between 50 and 500 bytes. This type of traffic carries time-sensitive data crucial to the AGV’s navigation, control, and safety, including information like position, speed, obstacle detection, and commands from the central control system. The predictable and regular intervals of NC traffic ensure reliable real-time performance.
@@
Video traffic is also periodic, reflecting its real-time data transmission requirements. Video is streamed from the AGV’s camera, supporting tasks such as object recognition, visual inspection, and remote monitoring. Packet sizes for video traffic vary from 1000 to 1500 bytes, corresponding to changes in scene complexity. The consistent periodicity of this traffic stream ensures stable video quality and minimizes latency. In contrast, BE traffic is irregular, with intervals ranging from 500 ms to 2 seconds and packet sizes between 30 and 1500 bytes. BE handles non-critical data, such as telemetry, diagnostics, or software updates, which are not essential for immediate AGV operations. This traffic often alternates between inactivity and bursts of high activity, making it suitable for low-priority background tasks where timing is flexible.
@@ IV Simulation Results and Discussion
Fig. 7 presents the results for the downlink scenario in an indoor factory environment, using different profiles (InF-SL, InF-DL, InF-SH, and InF-DH) and varying numbers of UEs in each case, with the distance defined as d2 across all configurations. Fig. 8 illustrates the corresponding results for the uplink scenario. These figures analyze the behaviour of the radio interface’s Hybrid Automatic Repeat reQuest (HARQ) mechanism under varying indoor factory conditions. The profiles explore the effects of sparse/dense clutter and low/high base station placement on transmission reliability. The HARQ error rate, measured as the ratio of failed transmissions to the total number of transmissions, serves as a key indicator of reliability for time-critical and non-critical data transmissions in the industrial environment. This comparison highlights how varying clutter, base station heights, and UE densities influence error performance under the given distance conditions.

### v2

@@ Abstract
While technologies such as Time-Sensitive Networking (TSN) improve deterministic behaviour, real-time functionality, and robustness of Ethernet, future industrial networks aim to be increasingly wireless. While wireless networks facilitate mobility, reduce cost, and simplify deployment, they do not always provide stringent latency constraints and highly dependable data transmission as required by many manufacturing systems. The advent of 5G, with its Ultra-Reliable Low-Latency Communication (URLLC) capabilities, offers potential for wireless industrial networks. 5G offers elevated data throughput, very low latency, and negligible jitter. As 5G networks typically include wired connections from the base station to the core network, integration of 5G with time-sensitive networking is essential to provide rigorous QoS standards. This paper assesses the scalability of 5G-TSN for various indoor factory applications and conditions using OMNET++ simulation. Our research shows that 5G-TSN has the potential to provide bounded delay for latency-sensitive applications in scalable indoor factory settings.
@@ I Introduction
With the rise of the Industrial Internet of Things (IIoT) and the push towards Industry 4.0, efforts are being made to enhance the flexibility, interoperability, and integration of industrial automation systems. The future of smart manufacturing will see the emergence of dynamic, reconfigurable production systems capable of rapid adaptation to evolving industry requirements. These environments will leverage technologies such as mobile sensing devices, robotic arms, automated guided vehicles (AGVs), and drones to facilitate the customization of products. Since traditional wired automation systems lack the adaptability needed for such changes, wireless technologies are becoming crucial to ensure real-time, deterministic connectivity for industrial processes.
Wireless technologies are advancing to support the demanding needs of industrial environments. Wireless TSN, building on the success of wired TSN, is emerging as a promising solution [1]. 5G and 802.11 are potential candidates for enabling wireless TSN. However, 5G, with its ultra-reliable, low-latency communication capabilities specifically designed for real-time communication through its URLLC profile, appears to be the current front-runner to become the wireless technology for Industry 4.0 applications.
5G and TSN integration promises to significantly advance real-time industrial communication. However, challenges such as precise time synchronization and efficient resource allocation remain. While early testbeds to study 5G-TSN performance are still under construction, further research is necessary to assess the scalability of 5G-TSN in industrial settings. While there are several studies on 5G-TSN systems, to the best of our knowledge, this is the first study to examine the scalability of 3GPP standardised 5G-TSN in a wireless manufacturing environment. The conducted simulation-based study provides a performance assessment of an industrial use case by integrating an indoor manufacturing model into a 5G-TSN industrial network model transmitting TSN traffic over 5G. Our research presents the following key contributions:
@@ II-A4 Resource management
Currently, 5G is recognized as the leading candidate for implementing wireless TSN solutions. In its Release 16, the 3rd Generation Partnership Project (3GPP) laid the groundwork for enabling mobile and wireless communications to achieve the speed and reliability required by Industry 4.0. This release introduces mechanisms to deliver ultra-reliable low-latency communication (URLLC) with sub-1 ms latency and highly precise synchronization, with an accuracy better than 1 $\mu$s, leveraging IEEE 802.1AS. Integration with IEEE 802.1 TSN is facilitated through the 5G Bridge framework, which incorporates critical elements such as the device-side TSN Translator (DS-TT), user equipment (UE), the radio access network (RAN), and the core network (5GC). The core network architecture is divided into user plane functions (UPFs), featuring the network-side TSN Translator (NW-TT), and control plane functions that include components for managing access, mobility (AMF), sessions (SMF), policies (PCF), and unified data (UDM) [2]. It is essential to understand the influence of wireless channels when incorporating TSN into 5G networks to ensure the effective management of TSN performance requirements. The next section delves into the unique characteristics of indoor factory environments and explores the use-case of 5G-TSN within this context.
@@ III-B Scenario/Use-Case
AGVs, originally developed for material handling in factories, have evolved into key mobile devices for flexible manufacturing. With 5G connectivity, they enhance operational efficiency by autonomously inspecting factories, improving safety, and streamlining logistics. Deploying multiple AGVs indoors optimizes workflow, making them a valuable investment for companies like Amazon, DHL, and Tesla. While real-world applications exist, researchers continue exploring scalability and implementation challenges for 5G-powered AGVs and AMRs. A notable example is a 5G-connected mobile industrial robot (MiR) equipped with a battery, LiDAR camera, video processing unit, and private 5G network access [4]. The setup integrates mobile edge computing for remote access and enhanced functionality, as shown in Fig. 2.
@@
In our 5G-TSN network scenario, traffic is categorized into three types. Automation Network Control (NC), Video - to partly support navigation, and Best Effort (BE). NC and Video are TSN traffic streams, with NC assigned the highest priority, Video a medium priority, and BE, a non-TSN stream, the lowest priority. NC traffic is periodic, where strict real-time behaviour is required. This type of traffic carries time-sensitive data crucial to the AGV’s navigation, control, and safety, including information like position, speed, obstacle detection, and commands from the central control system. Video traffic is periodic and is streamed from the AGV’s camera to support tasks such as object recognition, visual inspection, and remote monitoring. The consistent periodicity of this traffic stream ensures stable video quality and minimizes latency. BE handles non-critical data, such as telemetry, diagnostics, or software updates, which are not essential for immediate AGV operations. This traffic often alternates between inactivity and bursts of high activity, making it suitable for low-priority background tasks where timing is flexible.
To accurately model the behaviour of applications in our network, randomness is introduced in the initial production offset and start time of each traffic stream. The initial production offset determines when an application begins transmitting data after initialization, allowing for variability that simulates real-world delays in data generation. For instance, NC traffic, which requires strict real-time behaviour, has a smaller offset, enabling it to start transmission almost immediately. In contrast, Video and BE traffic experience longer offsets, with BE assigned the longest delay before transmission begins. Additionally, the start time, which dictates when the data transmission actually starts, varies for each application. NC traffic typically begins without delay, Video starts after a short delay, and BE starts with a more randomized time. This randomness in both the initial production offset and start time is crucial for ensuring that each traffic stream behaves realistically, reflecting the unpredictable nature of network traffic in real-world environments, where synchronization is rare, and application timing varies based on priority and type.
@@ IV Simulation Results and Discussion
Fig. 7 presents the results for the downlink scenario in an indoor factory environment, using different profiles (InF-SL, InF-DL, InF-SH, and InF-DH) and varying numbers of UEs in each case, with the distance defined as d2 across all configurations. Fig. 8 illustrates the corresponding results for the uplink scenario. These figures analyze the behaviour of the radio interface’s Hybrid Automatic Repeat reQuest (HARQ) mechanism under varying indoor factory conditions. The profiles explore the effects of sparse/dense clutter and low/high base station placement on transmission reliability. The HARQ error rate, measured as the ratio of failed transmissions to the total number of transmissions, serves as an indicator of reliability for time-critical and non-critical data transmissions. This comparison highlights how varying clutter, base station heights, and UE densities influence error performance under the given distance conditions.
@@ V Conclusion
Acknowledgments
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 13/RC/2077_P2 (CONNECT: the Research Ireland Centre for Future Networks).